Solve each system by the substitution method.

3x+y=2
-x-3y=6

Answer by checkley71(8403)

3X+Y=2 OR Y=-3X+2
-X-3(-3X+2)=6
-X+9X-6=6
8X=6+6
8X=12
X=12/8
X=1.5 ANSWER.
Y=-3*1.5+2
Y=-4.5+2
Y=-2.5 ANSWER.
PROOF
3*1.5-2.5=2
4.5-2.5=2
2=2